OpenStudy (anonymous):

An object initially at rest start to accelerate at a constant acceleration of 5 m/s2 at time t = 0 s. At time t = 3 s, the acceleration of the object reduces to 4 m/s2 and this continue until t = 10 s. a. At time t = 2 s, what is the speed of the object? b. At time t = 5 s, what is the speed of the object?

OpenStudy (anonymous):

object at time t=0 has zero velocity at t=1 due to the 5 meters PER SECOND acceleration v = 5m/s at t=2 v= 10m/s at t=3 v=15m/s at t = 4 v=19m/s at t=5 v=23m/s

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Thank you so much !